# Thunderscapes: Simulating the Dynamics of Mesoscale Convective System

### v1

@@
Abstract
A Mesoscale Convective System (MCS) is a collection of thunderstorms that function as a system, representing a widely discussed phenomenon in both the natural sciences and visual effects industries, and embodying the untamed forces of nature.In this paper, we present the first interactive, physically inspired mesoscale thunderstorms simulation model that integrates Grabowski-style cloud microphysics with atmospheric electrification processes. Our model simulates thunderclouds development and lightning flashes within a unified meteorological framework, providing a realistic and interactive approach for graphical applications. By incorporating key physical principles, it effectively links cloud formation, electrification, and lightning generation. The simulation also encompasses various thunderstorm types and their corresponding lightning activities.
1 INTRODUCTION
Thunderstorms represent the wild power of nature and are a common atmospheric element in visual effects (VFX) industry. Notable works, such as Horizon Forbidden West: Burning Shores and Ghost of Tsushima(see Figure LABEL:fig:thunderstorm_visuals), demonstrate the importance of realistic atmospheric effects. Modern applications demand tools for efficient and realistic simulation of thunderstorms. However, current general purpose VFX software, such as Houdini and Maya, lacks standardized toolkits specifically designed for creating atmospheric phenomena like thunderstorms. Artists often rely on traditional Eulerian fluid dynamics (e.g., stable fluids[Stam(1999)]) to simulate smoke and manually control the evolution of cloud shapes. Lightning, on the other hand, is typically generated using L-system based methods to create fractional geometries. However, these approaches do not account for the natural dynamics and interrelationship between thunderclouds and lightning.
Recent advances in computer graphics research have increasingly focused on cloud dynamics, incorporating atmospheric microphysical processes [Hädrich et al.(2020), Herrera et al.(2021), Amador Herrera et al.(2024)]. However, the critical role of atmospheric electrification in cloud development remains underexplored. This study aims to address this gap by integrating the electrification process, thus enhancing the consistency and realism of phenomena observed during thunderstorm events.
This paper presents an interactive, physically inspired method for simulating thunderstorm development. By coupling cloud microphysics with atmospheric hydrometeor electrification, our approach captures essential processes that contribute to the consistent formation of mesoscale thunderstorm phenomena.
@@ 1 INTRODUCTION
    We introduce a lightweight parameterization that enables users to interactively simulate various types of thunderstorms, along with their corresponding lightning effects.
@@
2 RELATED WORK
2.1 Simulating Thunderstorms in Computer Graphics
The simulation of atmospheric phenomena such as thunderstorms has been extensively explored through a variety of computational methods. Webanck et al. [Webanck et al.(2018)] proposed a procedural approach for generating cloudscapes, while Miyazaki et al. [Miyazaki et al.(2002)] simulated cumulus clouds by coupling computational fluid dynamics (CFD) with fundamental water transport equations. Such methods are also employed in Lagrangian-based particle methods. Ferreira et al. [Ferreira Barbosa et al.(2015)] and Zhang et al. [Zhang et al.(2020)] utilized position-based fluids (PBF) for adaptive cloud simulations. Smoothed Particle Hydrodynamics (SPH) techniques, as demonstrated by Goswami and Neyret [Goswami and Neyret(2017)], focus on real-time simulations of convective clouds. Additionally, Vimont et al. [Vimont et al.(2020)] proposed a hybrid, 2D-layered atmospheric model to simulate mesoscale skyscapes, contributing to a more detailed and interactive representation of large-scale cloud structures.
Some studies have specifically focused on the simulation of volcanic cloud dynamics. Lastic et al. [Lastic et al.(2022)] employed Lagrangian dynamics to simulate volcanic plumes and pyroclastic flows, while Pretorius et al. [Pretorius et al.(2024)] integrated volcanic eruptions with atmospheric simulations to produce coherent skyscapes. In their framework, lightning effects were incorporated, but the electrification process, which is essential for lightning formation, was not considered, resulting in a physically inaccurate simulation..
In the domain of lightning simulation, Reed and Wyvill [Reed and Wyvill(1994)], Kim and Lin [Kim and Lin(2007)], and Yun et al. [Yun et al.(2017)] developed methods for the efficient development of lightning branches, contributing to a more realistic representation of the method of electrical discharges.
Recently, more sophisticated microphysical schemes from atmospheric science have been incorporated into computer graphics research. Garcia-Dorado et al. [Garcia-Dorado et al.(2017)] and Hädrich et al. [Hädrich et al.(2020)] adopted the classic Kessler warm cloud microphysics scheme [Kessler(1969)] in cloud simulations, which has been instrumental in capturing the dynamics of cloud formation and precipitation. Herrera et al. [Herrera et al.(2021)] extended cloud simulations to include multiphase cloud dynamics, further enhancing the realism of weather phenomena. Amador Herrera et al. [Amador Herrera et al.(2024)] developed a framework to simulate hurricane and tornado dynamics, incorporating turbulent microphysics to model these extreme atmospheric events with greater accuracy.
2.2 Thunderstorm models in atmospheric sciences
@@
Kessler [Kessler(1969)] proposed a fundamental framework for the distribution and continuity of water substance in atmospheric circulations, which remains influential in the parameterization of the microphysics of warm clouds.
@@
One notable development in cloud microphysics is the work by Grabowski [Grabowski(1998)], who introduced an extended warm cloud microphysics scheme for large-scale tropical circulations. His method divides the parameterization of warm and cold clouds using a temperature interpolation scheme, a significant inspiration for our approach.
@@
In terms of thunderstorm electrification, Solomon et al. [Solomon et al.(2005)] introduced a 1.5-dimensional explicit microphysics thunderstorm model that incorporates a lightning parameterization, addressing key aspects of thunderstorm electrification. Furthermore, Mansell et al. [Mansell et al.(2002)] simulated three-dimensional branched lightning in a numerical thunderstorm model, providing insights on the complex activities of lightning formation. Barthe and Pinty [Barthe and Pinty(2007)] further advanced the field by simulating a supercell storm using a three-dimensional mesoscale model with an explicit lightning flash scheme, capturing the lightning activities specific to supercell thunderstorms.Mansell et al. [Mansell et al.(2010)] also examined the electrification of small thunderstorms using a two-moment bulk microphysics scheme, extending the understanding of lightning activity in small multicell thunderstorms.
Recent studies have focused on exploring the relationship between thundercloud structures and electrification processes. Formenton et al. [Formenton et al.(2013)] used a cloud electrification model to study the relationship between lightning activity and cloud microphysical structure, providing valuable insights into the interaction between cloud formation and lightning generation. More recently, Wu et al. [Wu et al.(2023)] employed coherent Doppler wind lidar to detect and analyze thundercloud structures, offering a novel method for observing thunderstorm activities.
These contributions highlight the ongoing efforts to simulate and understand thunderstorm activities, including cloud microphysics, lightning activity, and the complex interactions between these phenomena.
@@
Our work focuses on the interactive simulation of common thunderstorm types, based on the standard categories in Mesoscale Convective Systems (MCS). According to the National Severe Storms Laboratory (NSSL) [^1] https://www.nssl.noaa.gov/education/svrwx101/thunderstorms/types/, these types include single cell, multi-cell, squall line, and supercell thunderstorms. Additionally, our simulation includes common phenomena generated by thundercloud electrification, such as cloud-to-ground (CG) and intra-cloud (IC) lightning [^2] https://www.nssl.noaa.gov/education/svrwx101/lightning/types/.
@@
Figure 1: Schematic illustration of our thunderstorm microphysics scheme for hydrometeor phase transitions and the electrification process. The details will be further discussed in Chapter 4. (I) Blue frame: the transport of water substances during thundercloud development; (II)Orange frame: the accumulation of charge density due to the collision-coalescence of hydrometeors; (III) Red frame: the static charge, once it exceeds a given electric field threshold, triggers the lightning discharge process; (IV) Blue arrow: the neutralization of hydrometeors following the lightning flash.
3 OVERVIEW
@@
The primary motivation of our approach is to create consistent atmospheric phenomena during thunderstorms, particularly focusing on thundercloud development and dissipation, along with lightning flashes resulting from the thundercloud electrification processes,as illustrated in Figure 1. By simulating both the microphysics of thunderstorms and their electrodynamic properties, we aim to achieve realistic, interactive weather simulations suitable for graphical applications.
@@
Our model integrates key physical principles to simulate thunderstorm behavior at a mesoscale level. The thunderstorm microphysics component models atmospheric dynamics by coupling a Grabowski-style extended warm cloud microphysics scheme with hydrometeor electrification processing. This approach ensures that the microphysical processes driving cloud formation and growth are consistently linked to the electrification necessary for lightning generation. Additionally, we utilize the compact Poisson filter in our fluid dynamics algorithm to create an efficient simulation of the interactions between cloud particles and atmospheric forces. The compact Poisson filter provides an optimized balance between spectral and iterative methods, enabling high-performance simulations while maintaining stability and accuracy, which is essential for interactive graphics applications.
@@
The simulation is capable of modeling various thundercloud types, including single-cell (a simple, isolated thunderstorm cell), multicell (a cluster of thunderstorm cells with varying stages of development), squall line (a collection of storms that form a line hundreds of miles long, often characterized by a shelf cloud feature), supercell (a large, rotating thunderstorm with the potential to produce tornadoes), and atmospheric river (a narrow corridor of concentrated moisture that can fuel intense thunderstorms).
The lightning types simulated in our model include cloud-to-ground lightning (lightning that strikes from the cloud to the ground) and intra-cloud lightning (lightning occurring between different parts of the same cloud).For the simulation of lightning channel growth, we utilize the Dielectric Breakdown Model (DBM)[Kim and Lin(2007)], which captures the electrodynamics involved in the development of lightning.
The validation of our model is performed by comparing simulated results with real-world weather data. This includes cloud fraction profiles to assess cloud formation and structure, temperature, cloud coverage, and relative humidity evolution compared with real-time weather data from national weather services [^3] https://www.visualcrossing.com/weather/weather-data-services/, and lightning flash rate over time, evaluated against observed lightning activity.
Figure 2: Illustration of our thunderstorm microphysics scheme, which integrates the Grabowski-style extended warm cloud microphysics with the hydrometeor electrification process.
4 METHODOLOGY
@@ 4.1 Cloud Microphysics
$\frac{D\theta}{Dt}=\frac{L_{v}\theta}{c_{p}T}(COND-DIFF)$ (1)
$\frac{Dq_{v}}{Dt}=-COND+DIFF$ (2)
$\frac{Dq_{c}}{Dt}=COND-AUTC-ACCR$ (3)
$\frac{Dq_{p}}{Dt}=AUTC+ACCR-DIFF$ (4)
@@
$q_{vs}=\alpha q_{vw}+(1-\alpha)q_{vi}$ (5)
$q_{w}=\alpha q_{c},\,q_{i}=(1-\alpha)q_{c}$ (6)
$q_{r}=\alpha q_{p},\,q_{s}=(1-\alpha)q_{p}$ (7)
@@
$DIFF=DIFF_{r}+DIFF_{s}$ (8)
$AUTC=AUTC_{r}+AUTC_{s}$ (9)
$ACCR=ACCR_{r}+ACCR_{s}$ (10)
@@
$v_{p}=\alpha v_{r}+(1-\alpha)v_{s}$ (11)
The saturation vapor mixing ratio $q_{vs}$ combines contributions from vapor saturation over water and ice, weighted by $\alpha$[Yau and Rogers(1996)]:
$q_{vs}=\frac{380.16}{p}\left[\alpha\exp\left(\frac{17.67\cdot T}{T+243.50}\right)+(1-\alpha)\exp\left(\frac{24.46\cdot T}{T+272.62}\right)\right]$ (12)
@@
The diffusional growth rate $DIFF$ accounts for condensation and evaporation rates, integrating water and ice contributions [Dudhia(1989)],where $\beta_{r},\beta_{s}$ are phase-specific coefficients:
$DIFF=\beta_{r}\cdot\left(\alpha\cdot q_{vs}-q_{v}\right)^{+}+\beta_{s}\cdot\left((1-\alpha)\cdot q_{vs}-q_{v}\right)^{+}$ (13)
The autoconversion rate $AUTC$ describes cloud condensate aggregation into precipitation [Kessler(1995), Lin et al.(1983)]:
$AUTC=\beta_{r}\cdot(\alpha\cdot q_{c}-10^{-3})^{+}+\beta_{s}\cdot e^{0.025T}\cdot((1-\alpha)\cdot q_{c}-10^{-3})^{+}$ (14)
The accretion rate $ACCR$ models the collection of cloud condensate by precipitation particles [Morrison et al.(2015)]:
$ACCR=q_{c}\cdot q_{r}\cdot\left(\beta_{r}\cdot\alpha^{2}+\beta_{s}\cdot(1-\alpha)^{2}\right)$ (15)
4.2 Electrification
Based on the Reynolds-Brook theory of thunderstorm electrification [Latham and Miller(1965)], the fair-weather electric field ($E$) induces opposite charges on precipitation particles. Collisions between particles, influenced by their velocities and directions, result in partial charge neutralization and a residual net charge. Water droplet collisions generally coalesce without significant charge separation, while ice particle interactions are limited by brief contact times and low conductivity. Though water-ice and ice-riming collisions contribute, the inductive process alone is insufficient for thunderstorm-scale electric fields.This theory underpins the following mathematical model for electrification, with charge density ($\rho$) defined as:
$\rho=k\alpha(1-\alpha)\left(q_{c}^{2}v_{c}-q_{p}^{2}v_{p}\right)e^{-10q_{v}}$ (16)

### v3

@@
Abstract.
A Mesoscale Convective System (MCS) is a collection of thunderstorms that function as a system, representing a widely discussed phenomenon in both the natural sciences and visual effects industries, and embodying the untamed forces of nature.In this paper, we present the first efficient, physically based mesoscale thunderstorms simulation model that integrates Grabowski-style cloud microphysics with hydrometeor electrification processes. Our model simulates thunderclouds development and lightning flashes within a unified meteorological framework, providing a realistic and efficient approach for graphical applications. By incorporating key physical principles, it effectively links cloud formation, electrification, and lightning generation. The simulation also encompasses various thunderstorm types and their corresponding lightning activities. For more details, see the dynamic video: Thunderscapes Video.
Thunderstorm Modeling and Simulation,Fluid Simulation, Cloud Simulation,Lightning Simulation,Weather Simulation,Atmospheric Microphysics,Atmospheric Electrification
ccs: Computing methodologies Physical simulation
1. INTRODUCTION
Thunderstorms represent the wild power of nature and are a common atmospheric element in visual effects (VFX) industry. Notable works, such as Horizon Forbidden West: Burning Shores and Ghost of Tsushima(see Figure 2), demonstrate the importance of realistic atmospheric effects. Modern applications demand tools for efficient and realistic simulation of thunderstorms. However, current general purpose VFX software, such as Houdini and Maya, lacks standardized toolkits specifically designed for creating atmospheric phenomena like thunderstorms.
(a) Thunderstorm from *Ghost of Tsushima*. ©2020 Sony Interactive Entertainment LLC.
[…]
A scenic image from the video game Ghost of Tsushima showing dramatic landscapes and warrior scenes.
(b) Thunderstorm from *Horizon Forbidden West: Burning Shores*. ©2023 Sony Interactive Entertainment Europe.
[…]
A scene from the video game Horizon Forbidden West: Burning Shores, showcasing futuristic landscapes and action.
Figure 2. Comparison of atmospheric visual styles between *Ghost of Tsushima* and *Horizon Forbidden West: Burning Shores*.
Recent advances in computer graphics research have increasingly focused on cloud dynamics, incorporating atmospheric microphysical processes (Hädrich et al., 2020; Herrera et al., 2021; Amador Herrera et al., 2024). However, the critical role of atmospheric electrification in cloud development remains underexplored. This study aims to address this gap by integrating the electrification process, thus enhancing the consistency and realism of phenomena observed during thunderstorm events.
This paper presents an efficient, physically based method for simulating thunderstorm development. By coupling cloud microphysics with atmospheric hydrometeor electrification, our approach captures essential processes that contribute to the consistent formation of mesoscale thunderstorm phenomena.
@@ 1. INTRODUCTION
    We introduce a lightweight parameterization that enables users to efficiently simulate various types of thunderstorms, along with their corresponding lightning effects.
@@
2. RELATED WORK
2.1. Simulating Thunderstorms in Computer Graphics
The simulation of atmospheric phenomena such as thunderstorms has been extensively explored through a variety of computational methods. Webanck et al. (Webanck et al., 2018) proposed a procedural approach for generating cloudscapes, while Miyazaki et al. (Miyazaki et al., 2002) simulated cumulus clouds by coupling computational fluid dynamics (CFD) with fundamental water transport equations. Ferreira et al. (Ferreira Barbosa et al., 2015) and Zhang et al. (Zhang et al., 2020) utilized position-based fluids (PBF) for adaptive cloud simulations. Smoothed Particle Hydrodynamics (SPH) techniques, as demonstrated by Goswami and Neyret (Goswami and Neyret, 2017), focus on real-time simulations of convective clouds. Additionally, Vimont et al. (Vimont et al., 2020) proposed a hybrid, 2D-layered atmospheric model to simulate mesoscale skyscapes.
Some studies have specifically focused on the simulation of volcanic cloud dynamics. Lastic et al. (Lastic et al., 2022) employed Lagrangian dynamics to simulate volcanic plumes and pyroclastic flows, while Pretorius et al. (Pretorius et al., 2024) integrated volcanic eruptions with atmospheric simulations to produce coherent skyscapes.
In the domain of lightning simulation, Reed and Wyvill (Reed and Wyvill, 1994), Kim and Lin (Kim and Lin, 2007), and Yun et al. (Yun et al., 2017) developed methods for the efficient development of lightning branches, contributing to a more realistic representation of the method of electrical discharges.
Recently, more sophisticated microphysical schemes from atmospheric science have been incorporated into computer graphics research. Garcia-Dorado et al. (Garcia-Dorado et al., 2017) and Hädrich et al. (Hädrich et al., 2020) adopted the classic Kessler warm cloud microphysics scheme (Kessler, 1969) in cloud simulations. Herrera et al. (Herrera et al., 2021) extended cloud simulations to include multiphase cloud dynamics. Amador Herrera et al. (Amador Herrera et al., 2024) developed a framework to simulate hurricane and tornado dynamics.
2.2. Thunderstorm models in atmospheric sciences
@@
Kessler (Kessler, 1969) proposed a fundamental framework for the distribution and continuity of water substance in atmospheric circulations, which remains influential in the parameterization of the microphysics of warm clouds.
@@
One notable development in cloud microphysics is the work by Grabowski (Grabowski, 1998), who introduced an extended warm cloud microphysics scheme for large-scale tropical circulations. His method divides the parameterization of warm and cold clouds using a temperature interpolation scheme, a significant inspiration for our approach.
@@
In terms of thunderstorm electrification, Solomon et al. (Solomon et al., 2005) introduced a 1.5-dimensional explicit microphysics thunderstorm model that incorporates a lightning parameterization, addressing key aspects of thunderstorm electrification. Furthermore, Mansell et al. (Mansell et al., 2002) simulated three-dimensional branched lightning in a numerical thunderstorm model, providing insights on the complex activities of lightning formation. Barthe and Pinty (Barthe and Pinty, 2007) further advanced the field by simulating a supercell storm using a three-dimensional mesoscale model with an explicit lightning flash scheme, capturing the lightning activities specific to supercell thunderstorms.Mansell et al. (Mansell et al., 2010) also examined the electrification of small thunderstorms using a two-moment bulk microphysics scheme, extending the understanding of lightning activity in small multicell thunderstorms.
@@
Our work focuses on the efficient simulation of common thunderstorm types, based on the standard categories in Mesoscale Convective Systems (MCS). According to the National Severe Storms Laboratory (NSSL) [^1] https://www.nssl.noaa.gov/education/svrwx101/thunderstorms/types/, these types include single cell, multi-cell, squall line, and supercell thunderstorms. Additionally, our simulation includes common phenomena generated by thundercloud electrification, such as cloud-to-ground (CG) and intra-cloud (IC) lightning [^2] https://www.nssl.noaa.gov/education/svrwx101/lightning/types/.
@@
Figure 3. Schematic illustration of our thunderstorm microphysics scheme for hydrometeor phase transitions and the electrification process. The details will be further discussed in Chapter 4. (I) Blue frame: the transport of water substances during thundercloud development; (II)Orange frame: the accumulation of charge density due to the collision-coalescence of hydrometeors; (III) Red frame: the static charge, once it exceeds a given electric field threshold, triggers the lightning discharge process; (IV) Blue arrow: the neutralization of hydrometeors following the lightning flash.
[…]
A diagram illustrating the stages of hydrometeor phase transitions and electrification during thunderstorm development.
3. OVERVIEW
@@
The primary motivation of our approach is to create consistent atmospheric phenomena during thunderstorms, particularly focusing on thundercloud development and dissipation, along with lightning flashes resulting from the thundercloud electrification processes,as illustrated in Figure 3. By simulating both the microphysics of thunderstorms and their electrodynamic properties, we aim to achieve realistic, efficient weather simulations suitable for graphical applications.
@@
Our model integrates key physical principles to simulate thunderstorm behavior at a mesoscale level. The thunderstorm microphysics component models atmospheric dynamics by coupling a Grabowski-style extended warm cloud microphysics scheme with hydrometeor electrification processing. This approach ensures that the microphysical processes driving cloud formation and growth are consistently linked to the electrification necessary for lightning generation.
@@
The validation of our model involves presenting spatially simulated thundercloud structures alongside meteorological characteristics and comparing the temporally simulated results with real-world weather data. This includes analyzing cloud fraction profiles to assess cloud formation and structure, tracking the evolution of cloud coverage against real-time data from national weather services [^3] https://www.visualcrossing.com/weather/weather-data-services/, and evaluating the temporal variation in lightning flash rates.
Figure 4. Illustration of our thunderstorm microphysics scheme, which integrates the Grabowski-style extended warm cloud microphysics with the hydrometeor electrification process.
[…]
A schematic representation of the thunderstorm microphysics scheme, coupling the extended warm cloud microphysics with the electrification process of hydrometeors.
4. METHODOLOGY
@@ 4.1. Cloud Microphysics
(1) $\frac{D\theta}{Dt}=\frac{L_{v}\theta}{c_{p}T}(COND-DIFF)$
(2) $\frac{Dq_{v}}{Dt}=-COND+DIFF$
(3) $\frac{Dq_{c}}{Dt}=COND-AUTC-ACCR$
(4) $\frac{Dq_{p}}{Dt}=AUTC+ACCR-DIFF$
@@
(5) $q_{vs}=\alpha q_{vw}+(1-\alpha)q_{vi}$
(6) $q_{w}=\alpha q_{c},\,q_{i}=(1-\alpha)q_{c}$
(7) $q_{r}=\alpha q_{p},\,q_{s}=(1-\alpha)q_{p}$
@@
(8) $DIFF=DIFF_{r}+DIFF_{s}$
(9) $AUTC=AUTC_{r}+AUTC_{s}$
(10) $ACCR=ACCR_{r}+ACCR_{s}$
@@
(11) $v_{p}=\alpha v_{r}+(1-\alpha)v_{s}$
The saturation vapor mixing ratio $q_{vs}$ combines contributions from vapor saturation over water and ice, weighted by $\alpha$(Yau and Rogers, 1996):
(12) $q_{vs}=\frac{380.16}{p}\left[\alpha\exp\left(\frac{17.67\cdot T}{T+243.50}\right)+(1-\alpha)\exp\left(\frac{24.46\cdot T}{T+272.62}\right)\right]$
@@
The diffusional growth rate $DIFF$ accounts for condensation and evaporation rates, integrating water and ice contributions (Dudhia, 1989),where $\beta_{r},\beta_{s}$ are phase-specific coefficients:
(13) $DIFF=\beta_{r}\cdot\left(\alpha\cdot q_{vs}-q_{v}\right)^{+}+\beta_{s}\cdot\left((1-\alpha)\cdot q_{vs}-q_{v}\right)^{+}$
The autoconversion rate $AUTC$ describes cloud condensate aggregation into precipitation (Kessler, 1995; Lin et al., 1983):
(14) $AUTC=\beta_{r}\cdot(\alpha\cdot q_{c}-10^{-3})^{+}+\beta_{s}\cdot e^{0.025T}\cdot((1-\alpha)\cdot q_{c}-10^{-3})^{+}$
The accretion rate $ACCR$ models the collection of cloud condensate by precipitation particles (Morrison et al., 2015):
(15) $ACCR=q_{c}\cdot q_{r}\cdot\left(\beta_{r}\cdot\alpha^{2}+\beta_{s}\cdot(1-\alpha)^{2}\right)$
4.2. Electrification
Based on the Reynolds-Brook theory of thunderstorm electrification (Latham and Miller, 1965), the fair-weather electric field ($E$) induces opposite charges on precipitation particles. Collisions between particles, influenced by their velocities and directions, result in partial charge neutralization and a residual net charge. This theory underpins the following mathematical model for electrification, with charge density ($\rho$) defined as:
(16) $\rho=k\alpha(1-\alpha)\left(q_{c}^{2}v_{c}-q_{p}^{2}v_{p}\right)e^{-10q_{v}}$